# Geminga: A Window of the Role Played by Local Halo in the Cosmic Ray Propagation Process

### v1

@@ 3.3 Contribution of the Geminga to the positron spectrum
Geminga is a middle-aged pulsar located about from Earth, with an age of and a spin-down power of (Faherty et al., 2007). Its distance and age provide strong grounds to believe it could be a major source of the positron excess. The spectrum of positrons obtained near Earth, as shown in Figure 4, is constrained by the data from Fermi, HAWC, and LHAASO. While the spectrum is nearly consistent with the observational data, it is difficult to distinguish between electrons and positrons in the halo, so Geminga may contribute less than of the observed positrons. This suggests that other nearby candidate sources or alternative physical processes, such as dark matter annihilation or proton-proton (p-p) interactions of cosmic rays from nearby sources, could also play a significant role and should not be ignored.
@@ 4 Conclusion and Summary
This paper investigates the transport and radiation properties of electrons/positrons in the Geminga halo using observational data from LHAASO, Fermi, and HAWC, with the GALPROP package as the simulation tool. It expands on the role of pulsar halos in the cosmic ray propagation process, explores the spatial dependence of diffuse gamma radiation, and tests the spatially dependent cosmic ray propagation models.
We reproduce the non-thermal radiation energy spectrum from the Geminga halo, and the extended radiation maps in the 1-10 GeV and 1-100 TeV energy ranges. Our findings suggest that only a tiny fraction of the low-energy electrons/positrons injected by the Geminga pulsar may escape from its surrounding slow diffusion region and produce radiation in the extended area. The high-energy radiation phenomena near the pulsar (including the extended radiation observed by HAWC, which spans more than 2.6 degrees, and the “bump” structure in the diffuse gamma-ray radiation distribution observed by LHAASO in the region of the pulsar’s silver disk) are primarily caused by the contribution of high-energy electrons/positrons that escape the slow diffusion region. Our calculation results strongly support the idea that local sources mainly contribute to high-energy cosmic ray phenomena. On the one hand, the local cosmic ray halo prevents the propagation of external cosmic rays, and on the other hand, it limits the escape of local cosmic rays. Furthermore, diffuse gamma-ray radiation in the Milky Way shows spatial dependence. We believe that LHAASO will be able to observe the extended radiation from the Geminga halo and other TeV halos in the future, and more precise observations will further validate our conclusions.
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (12275279,12375103).
@@
where represents the CR density per unit of total particle momentum at position , describes the source term, denotes the spatial diffusion coefficient, is the convection velocity and and are the timescales for loss by fragmentation and radioactive decay, respectively. The CR diffusion depends on the distribution of CR sources , and the diffusion coefficient is described as(Guo et al., 2016; Guo & Yuan, 2018)
@@
where is the Alfvén speed, and w is the ratio of magneto-hydrodynamic wave energy density to the magnetic field energy density, which can be fixed to 1.

### v2

@@ 3.3 Contribution of the Geminga to the positron spectrum
Geminga is a middle-aged pulsar located about from Earth, with an age of and a spin-down power of (Faherty et al., 2007). Its distance and age provide strong grounds to believe it could be a major source of the positron excess. The spectrum of positrons obtained near Earth, as shown in Figure 8, is constrained by the data from Fermi, HAWC, and LHAASO. While the spectrum is nearly consistent with the observational data, it is difficult to distinguish between electrons and positrons in the halo, so Geminga may contribute less than of the observed positrons. Besides, the Geminga contributes to the positron flux at Earth strongly depend on the chosen injection slope of the spectrum and size of SDZ(Di Mauro et al., 2019; Jóhannesson et al., 2019). The softer injection spectra have more low-energy electrons, and there is more confinement for low-energy electrons for larger sizes of the SDZ. A smaller SDZ leads to a larger flux of positrons at Earth because of the larger effective diffusion coefficient, and larger values of also result in a larger positron flux at Earth in the observed energy range. In addition, it may be suggested that other nearby candidate sources or alternative physical processes, such as dark matter annihilation or proton-proton (p-p) interactions of cosmic rays from nearby sources, could also play a significant role and should not be ignored.
@@ 4 Conclusion and Summary
This paper investigates the transport and radiation properties of electrons/positrons in the Geminga halo using observational data from LHAASO, Fermi-LAT, and HAWC, with the GALPROP package as the simulation tool. It expands on the role of pulsar halos in the cosmic ray propagation process, explores the spatial dependence of diffuse gamma radiation, and tests the physical framework of the spatially dependent cosmic ray propagation.
The previous findings suggest that Fermi-LAT observations reveal no significant extended emission within the region surrounding Geminga in the energy range of 10 to 100 GeV (Xi et al., 2019), while the corresponding emission is detected in a larger field (Di Mauro et al., 2019). We reproduce the non-thermal radiation energy spectrum from the Geminga halo, and simulate the extended radiation maps in the 1-10 GeV and 1-100 TeV energy ranges. Our calculated results support that this may result from the pulsar’s proper motion combined with its birth in a slow-diffusion region.
We find that the signal leakage from the local Geminga halo may contribute to some extent to the fluctuating structure of the diffuse gamma-ray radiation observed by LHAASO. Furthermore, such leakage from local sources may contribute to the diffuse gamma-ray “excess”. However, the contributions from signal leakage and local gas uncertainties cannot be disentangled, leaving it uncertain which effect dominates the measured fluctuations. We hope that LHAASO will be able to observe the extended radiation from the Geminga halo and other TeV halos in the future, and more precise observations will further study and disentangle signal leakage and other effects.
@@ Acknowledgements
We thank Silvia Manconi for providing the XXM data. This work is supported in China by National Key RD program of China under the grant 2024YFA1611402, and supported by the National Natural Science Foundation of China (12333006, 12275279, 12375103).
@@
where represents the CR density per unit of total particle momentum at position , describes the source term, denotes the spatial diffusion coefficient, is the convection velocity and and are the timescales for loss by fragmentation and radioactive decay, respectively. However, this work does not account for Galactic convection, so the convection velocity is set to zero. The CR diffusion depends on the distribution of CR sources , and the diffusion coefficient is described as(Guo et al., 2016; Guo & Yuan, 2018)
@@
where is the Alfvén speed, and is the ratio of magneto-hydrodynamic wave energy density to the magnetic field energy density, which can be fixed to 1. Figures 9 and 10 present the comparison between the proton spectra and the B/C ratio obtained from observations and the predictions of the SDP model, respectively.